Q = -L^3 + 5L^2 + 10L

Derrive MPL and APL

At what level of labor employment does the total production begin to decline ?


1
Expert's answer
2022-03-08T12:39:51-0500

MPL=(Q)'=-3L2+10L+10

APL=Q/L=-L2+5L+10

the total production begin to decline when MP=0

-3L2+10L+10=0

3L2-10L-10=0

D=(-10)2-4*(-10)*3=100+120=220

L=(10+14.83)/6=4.14